Good morning pls help me with this situation. :)
"Freddie is five times as old as Elena, but in four years, he will be 8 yrs more than thrice as old as her. Find the age of each now."
Thank you pls reply back! :>

Answer by ankor@dixie-net.com(22740)

Good morning pls help me with this situation. :)
"Freddie is five times as old as Elena, but in four years, he will be 8 yrs more than thrice as old as her. Find the age of each now."
:
let f = Fred's present age
let e = Elena's age
:
Write an equation for each statement
"Freddie is five times as old as Elena,"
f = 5e
" but in four years, he will be 8 yrs more than thrice as old as her."
Add 4 to both ages.
f + 4 = 3(e+4) + 8
f + 4 = 3e + 12 + 8
f + 4 = 3e + 20
f = 3e + 20 - 4
f = 3e + 16
Replace f with 5e from the 1st statement
5e = 3e + 16
5e - 3e = 16
2e = 16
e = 8 yrs is E's age
find f
f = 5(8)
f = 40 yrs if F's age
:
:
Check solutions in the 2nd statememt
"in four years, he will be 8 yrs more than thrice as old as her."
40 + 4 = 3(8 + 4) + 8
44 = 3(12) + 8
:
:
You can use this method to solve all these type of problems